Solve. Where appropriate, include approximations to the nearest thousandth. If no solution exists, state this.
In (2x+3)=2

Answer by edjones(7569)

In (2x+3)=2
e^(In (2x+3))=e^2
2x+3=e^2
2x=e^2-3
x=(e^2-3)/2
=2.195
.
Ed